A new cruise ship line has just launched 3 new​ ships: the Pacific​ Paradise, the Caribbean​ Paradise, and the Mediterranean Paradise. The Caribbean Paradise has 30 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 35 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 1165.
Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52776)

.
This problem was  SO MANY TIMES  posted to  AND SOLVED at  the forum
that I prepared  SPECIAL LESSON  for you

    - More solved word problems on a single linear equation,   Problem 2

in this site.  Go  to there  and learn  from there.


                    H a p p y   l e a r n i n g  ! !


https://www.algebra.com/algebra/homework/word/evaluation/Evaluation_Word_Problems.faq.question.363881.html
https://www.algebra.com/algebra/homework/word/misc/Miscellaneous_Word_Problems.faq.question.834312.html
http://www.algebra.com/algebra/homework/word/evaluation/Evaluation_Word_Problems.faq.question.830722.html
http://www.algebra.com/algebra/homework/equations/Equations.faq.question.789864.html

=============

The major lesson and the major thing which you need to learn on and to know about it  is that this problem is for one single equation in one unknown.


Answer by stanbon(75887)

A new cruise ship line has just launched 3 new​ ships: the Pacific​ Paradise, the Caribbean​ Paradise, and the Mediterranean Paradise.
The Caribbean Paradise has 30 more deluxe state rooms than the Pacific Paradise. The Mediterranean Paradise has 35 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise.
Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 1165.
-----
Equations:
P + C + M = 1165
C = P + 30
M = 4P-35
--------------------
Substitute and solve for "P":
P + P+30 + 4P-35 = 1165
6P = 1170
P = 195 (# on PP)
C = 195+30 = 225 (# on CP)
M = 4P-35 = 745 (# on MP)
------------
Cheers,
Stan H.
-----------